OpenStudy (anonymous):

Please help !! Suppose a computer virus begins by infecting 8 computers in the first hour after it is released. Each hour after that, each newly infected computer causes 8 more computers to become infected. The function y = 8x models this situation. Make a table with integer values of x from 1 to 4. Could someone please help me with this,, i do not understand it.

OpenStudy (mysesshou):

It sounds like you substitute 1 through 4 for x into the equation. Then make a table using y and x. Hopefully someone will double check me. :) So, if x=1, then y=8x = 8*(1) = 8 So, x=1, y=8 If x=2, then y= 8x = 8*(2) = 16 x=2, y=16 if x=3, y = 8*(3) = 24 x=3, y=24 if x=4, y= 8*(4) = 32 x=4, y=32 put these into a table: x y 1 8 2 16 3 24 4 32 where the x and y are underlined to show they are the headers of the table. hope this helps some :)

OpenStudy (anonymous):

thanks so much!

OpenStudy (mysesshou):

welcome! hope it makes sense ok. :)

OpenStudy (anonymous):

yes it does! :)